In a family there are two cars. The sum of the average miles per gallon obtained by the two cars in a particular week is 65. The first car has consumed 25 gallons during that week, and the second has consumed 35 gallons, for a total of 2025 miles driven by the two cars combined. What was the average gas mileage obtained by each of the two cars in that week?
Answer by mananth(16946)

In a family there are two cars. The sum of the average miles per gallon obtained by the two cars in a particular week is 65. The first car has consumed 25 gallons during that week, and the second has consumed 35 gallons, for a total of 2025 miles driven by the two cars combined. What was the average gas mileage obtained by each of the two cars in that week?
Let the first car travel x miles in the week
Average of first car = x/25
The second car travesl 2025-x miles in that week
The average of second car = 2025-x / 35
x/25 + 2025-x / 35 = 65
35x+25*2025 -25x / 35*25* 65
10x+50625 = 56875
10x= 56875-50625
10x=6250

x= 625 miles distance traveled by first car
2025-625= 1400 miles second car
Ananth